# Representation Learning with Conditional Information Flow Maximization

### v1

@@ Abstract
This paper proposes an information-theoretic representation learning framework, named conditional information flow maximization, to extract noise-invariant sufficient representations for the input data and target task. It promotes the learned representations have good feature uniformity and sufficient predictive ability, which can enhance the generalization of pre-trained language models (PLMs) for the target task. Firstly, an information flow maximization principle is proposed to learn more sufficient representations by simultaneously maximizing both input-representation and representation-label mutual information. In contrast to information bottleneck, we handle the input-representation information in an opposite way to avoid the over-compression issue of latent representations. Besides, to mitigate the negative effect of potential redundant features, a conditional information minimization principle is designed to eliminate negative redundant features while preserve noise-invariant features from the input. Experiments on 13 language understanding benchmarks demonstrate that our method effectively improves the performance of PLMs for classification and regression. Extensive experiments show that the learned representations are more sufficient, robust and transferable.
@@ 1 Introduction
The goal of deep representation learning LeCun et al. (2015) is to transform the raw observational data into low-dimensional representations that are essential for various downstream tasks. In recent years, information-theoretic representation learning has been widely studied, aiming to discover useful representations in a principled manner. The InfoMax principle Linsker (1988) has extensive applications in the field of self-supervised representation learning van den Oord et al. (2018); Hjelm et al. (2019); Tschannen et al. (2020). In supervised scenarios, minimizing the standard cross-entropy is actually equivalent to maximizing the mutual information between the representations and the target task Boudiaf et al. (2020); Amjad and Geiger (2020). But InfoMax tends to preserve potential redundant features that are irrelevant to the given target, leading to biased representations. Another noteworthy line of information-theoretic research Alemi et al. (2017); Belghazi et al. (2018); Fischer (2020); An et al. (2023); Hu et al. (2024), built upon the information bottleneck (IB) principle Tishby et al. (1999); Tishby and Zaslavsky (2015), aims to discover compact and informative representations that can reduce redundant features from the inputs. IB seeks to find a maximally compressed representation of the input that preserves as much information as possible about the target, striking a balance between compression and prediction.
However, in the information flow Goldfeld et al. (2019) of neural networks, directly reducing the mutual information between the input $X$ and representations $Z$ would violate the sufficiency constraint, and may lose the necessary information for the target task $Y$. Under the Markov chain constraint $Y\rightarrow X\rightarrow T$, it’s hard to determine beforehand how close we are to optimal compression, and this can easily lead to the over-compression issue of latent representations Fischer (2020); Hu et al. (2024). As a result, current IB-based methods would yield insufficient representations for the target task, and hamper prediction ability of neural networks..
@@
Firstly, we propose an information flow maximization (IFM) principle to learn more sufficient representations for the input and target. It simultaneously maximizes both input-representation and representation-label mutual information. Maximizing input-representation information $I(X;Z)$ ensures sufficiency of the representations for the input $X$ and preserves information relevant to the target $Y$, and maximizing representation-label information $I(Y;Z)$ captures necessary information relevant to the target $Y$. In this way, the learned representation can be of good uniformity and sufficient predictive ability. Unlike IB that minimizes input-representation information, we handle the information in an opposite way to avoid the over-compression issue of latent representations.
@@ 2.1 Information Flow Maximization
As shown in Figure 1, regions being minimized by $I(X;Z)$ overlap with relevancy regions being maximized by $I(Z;Y)$. To preserve informative signals from the input $X$ relevant to the target $Y$ (i.e., region ①) as much as possible in the information flow of $X\rightarrow Z$, we use the idea of InfoMax to maximize $I(X;Z)$. Besides, we also maximize $I(Y;Z)$ to capture necessary information relevant to the target task $Y$. Combing the above two terms, the principle of information flow maximization (IFM) simultaneously maximizes both input-representation and representation-label mutual information. It can be formulated as the maximization of the following Lagrangian,
@@
In Equation (1), the second term that maximizing input-representation information $I(Y;Z)$ ensures sufficiency of the representations for the input $X$ and preserves informative signals from $X$ relevant to the target $Y$. The first term that maximizing representation-label information $I(Y;Z)$ captures necessary information relevant to the target $Y$. In this way, the learned representation by IFM can be of good uniformity and sufficient predictive ability. Unlike the IB principle that minimizes input-representation information, we handle the information in an opposite way to avoid the over-compression issue of latent representations.
@@
The implementation of IFM contains two vital terms, i.e., maximizing $I(Y;Z)$ and $I(X;Z)$. First, we maximize the lower bound of $I(Y;Z)$ by estimating the conditional entropy of the target $Y_{t}$ given representations $Z_{t}$. Following Amjad and Geiger (2020) and Hu et al. (2024), we use cross-entropy (CE) and mean squared error (MSE) as the estimators of maximizing $I(Y;Z)$ for classification and regression tasks, respectively. Then, maximizing $I(X;Z)$ can be optimized by using the mutual information estimators, e.g., InfoNCE [^2] InfoNCE is a softmax-based version of noise-contrastive estimation Gutmann and Hyvärinen (2010). van den Oord et al. (2018) and MINE Belghazi et al. (2018).
Here, we take InfoNCE as the default estimator for maximizing $I(X;Z)$. According to the information flow of $Z\leftarrow X\rightarrow Z^{\prime}$, the Markov chain rule states that $I(X;Z)\geq I(Z;Z^{\prime})$. Maximizing the lower bound of $I(X;Z)$, i.e., $I(Z;Z^{\prime})$, can be optimized by using InfoNCE estimator. Specifically, for a sample $x_{i}$ with the representation $z_{i}$, its positive key is the augmented sample obtained by dropout, and its negative keys are the other samples in the batch $N$. In this case, we have,
@@
where $z_{i}^{+}$ is the positive key, which is the augmented sample obtained by dropout. $z_{j}^{-}$ represents the negative keys, which are the representations of the other samples in the batch. $K$ is the number of negative keys, i.e., $22N-$ 2. $\tau>0$ is a scalar temperature parameter that controls the concentration or separation of probability distribution.
Alternatively, we also present how to utilize the MINE (mutual information neural estimator) estimator to maximize $I(X;Z)$. Specifically, we use a neural network to approximate the lower bound of $I(X;Z)$, i.e.,
@@
where ${\mathbb{P}_{{X}{Z}}}$ is the joint distribution and ${\mathbb{P}_{X}\otimes\mathbb{P}_{Z}}$ is the product of the marginals. $T_{\omega}$ is a function parametrized by an MLP (a fully-connected neural network with three hidden layers) with the parameter $\omega\in\Omega$. The expectations of $I_{\Omega}(X;Z)$ are estimated by shuffling the samples from the joint distribution along the batch axis. Additionally, an exponential moving average is used to make a biased estimate.
@@ 2.2 Conditional Information Minimization
Inspired by Terzi et al. (2021), we use adversarial training Goodfellow et al. (2015); Miyato et al. (2017) to approximately estimate the minimization of the Fisher information $F_{z|x,y}$ given the input $X$ and target $Y$. Formally, denote $(x,y)$ as a mini-batch input sampled from distribution $D$ and $p(y|x;{\theta})$ as a model with the parameter $\theta$. The worst-case perturbation $\delta$ under the given $Y$ can be computed by the backpropagation in the network, i.e., $\max_{\|\delta\|_{q}\leq\epsilon}{{KL}(p(\cdot|x;{\theta_{\delta}});p(\cdot|x;{\theta}))}$ with a Lagrange constraint $\mathcal{L}_{IFM}(x,y;\theta_{\delta})$ and an $L_{q}$ norm constraint. At each step of training, we identify the adversarial perturbations $\delta$ against the current network with the parameter $\hat{\theta}$, and put them on the weights of first several hidden layers of the network. When using a Transformer-based network (e.g., RoBERTa) as the model backbone, we probabilistically put the perturbations on the weights of an embedding layer and the first hidden layer of the Transformer based on a predefined perturbation rate. With a linear approximation Goodfellow et al. (2015), an $L_{q}$ norm-ball constraint, a certain radius $\epsilon$ for $\delta$, and the original objective $\mathcal{L}_{IFM}$, the formulation of adversarial objective $\mathcal{L}_{CIM}$ is,
@@ 3.3 Ablation Study
Table 5: Transferability results (%) of GoEmotions $\rightarrow$ Others and SentiEval $\rightarrow$ Others with RoBERTa backbone. GoEmotions and SentiEval are the source datasets for pre-training. Pre-train Obj. indicates the pre-training objective. Cls. refers to the classifier when fine-tuning. Detailed results of SentiEval $\rightarrow$ Others are listed in Appendix C.2.
@@ 3.4 Generalization Evaluation
We choose emotion-related benchmarks (e.g., EmotionEval, ISEAR, MELD, and GoEmotions), which are collected from different domains and aim to predict the label of different emotion taxonomies. To implement OOD scenarios, we train the model on the original training set from a source domain, select the best model based on the validation set of the source domain, and test on the test set of a target domain. To avoid the interference of label mapping bias between different taxonomies, each model is trained on the dataset with coarse-grained taxonomy to predict the label for another dataset with fine-grained taxonomy (i.e., GoEmotions). Table 4 shows the performance under OOD scenarios. CIFM obtains the best results on all OOD settings. Comparing to CE, CIFM achieves +7.9% improvements in terms of average scores. The results exhibit that CIFM can promote the learning of more sufficient representations for target task than others, which enhances the generalization in handling OOD scenarios across different domain shifts.
@@ 3.7 Representation Quality Evaluation
To evaluate the quality of the learned representations, we use two metrics to measure the sufficiency of both the input and target task. Following Wang and Isola (2020), we choose uniformity metric to measure the preserved maximal information of hidden representation $Z$ from the input. Following Hu et al. (2023), we apply adjusted rand index (ARI) metric to assess the preserved maximal information of output representation $Y^{\prime}$ for label structure. The lower the uniformity, the better sufficiency relevant to the data. Conversely, a higher ARI indicates a better sufficiency relevant to the task.
@@ 4 Conclusion
We propose an information-theoretic CIFM framework to learn sufficient representations for the input and target, as well as eliminates negative redundant features. Firstly, an IFM principle is proposed to learn more sufficient representations by simultaneously maximizing both input-representation and representation-label mutual information. Besides, a CIM principle is designed to eliminate negative redundant features while preserve noise-invariant features from the input. Experiments on 13 language understanding benchmarks demonstrate that CIFM effectively improves the performance of PLMs for classification and regression. Extensive experiments show that the learned representations are more sufficient, robust and transferable.
@@
This paper introduces a new information-theoretic framework to extract noise-invariant sufficient representations for the input data and target task. It promotes learned representations with good uniformity and sufficient predictive ability, thereby enhancing the generalization of pre-trained language models for the target task. However, the technique is studied only for different classification and regression tasks; its effectiveness in generation tasks remains to be further validated. Additionally, the principles proposed in this paper require more in-depth theoretical investigations.
@@ Appendix A Related Work
Information-theoretic representation learning has become the workhorse of several impressive deep learning achievements, ranging from practical applications Pereyra et al. (2017); Alemi et al. (2017); Belghazi et al. (2018) to theoretical investigations Xu and Raginsky (2017); Steinke and Zakynthinou (2020); Kawaguchi et al. (2023). In the context of the training principle, information-theoretic representation learning approaches can be broadly categorized into InfoMax-based and IB-based methods.
@@
In the field of machine learning, the information maximization (InfoMax) principle was first introduced by Linsker (1988). It suggests that a linear or nonlinear network can be viewed as an information channel and the goal is to transmit as much information as possible from the input data through the network. Recently, InfoMax has extensive applications in the field of self-supervised representation learning van den Oord et al. (2018); Hjelm et al. (2019); Tschannen et al. (2020). The objective is to maximize the mutual information (MI) between the input and the latent representations. In supervised scenarios, minimizing the standard cross-entropy is actually equivalent to maximizing the mutual information between the representations and the target task Boudiaf et al. (2020); Amjad and Geiger (2020). Moreover, Pereyra et al. (2017) add the negative entropy to the negative log-likelihood to penalize overconfident output distributions in supervised learning.
@@
The information bottleneck (IB) principle Tishby et al. (1999); Tishby and Zaslavsky (2015) aims to balance the trade-off between the compression of representation and the power of predicting. Alemi et al. (2017) propose an efficient variational estimation method based on the IB principle. Peng et al. (2019) use IB to constrain information flow in the discriminator for better adversarial learning. Then, Mahabadi et al. (2021) use VIB to suppress irrelevant features when fine-tuning on low-resource language understanding tasks. Due to the reliance on a variational approximation, VIB require a tractable density for the approximate posterior. For tractable application of IB in a continuous setting, Belghazi et al. (2018) propose a mutual information neural estimation method with IB principle, named MINE-IB in this paper. And Ragonesi et al. (2021) employ MINE-IB to learn unbiased representations. Fischer (2020) and Ramé and Cord (2021) introduce a conditional mutual information term to alleviate the over- and under-compression issue of IB. An et al. (2023) explicitly use the conditional entropy of the stochastic embedding as a confidence indicator and encourage the model to assign larger variance to more certain inputs. Recently, Hu et al. (2024) propose a probabilistic coding framework to mitigate the over-compression issue of IB by simultaneously performing compression and prediction using a shared learnable network. They also introduced a regularization term to promote class-level uniformity for improved task prediction.

### v2

@@ Abstract
This paper proposes an information-theoretic representation learning framework, named conditional information flow maximization, to extract noise-invariant sufficient representations for the input data and target task. It promotes the learned representations have good feature uniformity and sufficient predictive ability, which can enhance the generalization of pre-trained language models (PLMs) for the target task. Firstly, an information flow maximization principle is proposed to learn more sufficient representations for the input and target by simultaneously maximizing both input-representation and representation-label mutual information. Unlike the information bottleneck, we handle the input-representation information in an opposite way to avoid the over-compression issue of latent representations. Besides, to mitigate the negative effect of potential redundant features from the input, we design a conditional information minimization principle to eliminate negative redundant features while preserve noise-invariant features. Experiments on 13 language understanding benchmarks demonstrate that our method effectively improves the performance of PLMs for classification and regression. Extensive experiments show that the learned representations are more sufficient, robust and transferable.
@@ 1 Introduction
The goal of deep representation learning LeCun et al. (2015) is to transform the raw observational data into low-dimensional representations that are essential for various downstream tasks. In recent years, information-theoretic representation learning has been widely studied, aiming to discover useful representations in a principled manner. The InfoMax principle Linsker (1988) has extensive applications in the field of self-supervised representation learning van den Oord et al. (2018); Hjelm et al. (2019); Tschannen et al. (2020). In supervised scenarios, minimizing the standard cross-entropy is actually equivalent to maximizing the mutual information between the representations and the target task Achille and Soatto (2018); Boudiaf et al. (2020). But InfoMax tends to preserve potential redundant features that are irrelevant to the given target, leading to biased representations. Another noteworthy line of information-theoretic research is built upon the information bottleneck (IB) principle Tishby et al. (1999); Tishby and Zaslavsky (2015), which aims to discover compact and informative representations that can reduce redundant features from the inputs Alemi et al. (2017); Belghazi et al. (2018); Fischer (2020); An et al. (2023); Hu et al. (2024). IB seeks to find a maximally compressed representation of the input that preserves as much information as possible about the target, striking a balance between compression and prediction.
However, in the information flow Goldfeld et al. (2019) of neural networks, directly reducing the mutual information between the input $X$ and representations $Z$ would violate the sufficiency constraint, and may lose the necessary information for the target task $Y$. Under the Markov chain constraint $Y\rightarrow X\rightarrow Z$, it’s hard to determine beforehand how close we are to optimal compression, and this can easily lead to the over-compression issue of latent representations Fischer (2020); Hu et al. (2024). As a result, current IB-based methods would yield insufficient representations for the target task, and hamper prediction ability of neural networks.
@@
Firstly, we propose an information flow maximization (IFM) principle to learn more sufficient representations for the input and target. It simultaneously maximizes both input-representation and representation-label mutual information. Maximizing input-representation information $I(X;Z)$ ensures sufficiency of the representations for the input $X$ and preserves information relevant to the target $Y$, and maximizing representation-label information $I(Y;Z)$ captures necessary information relevant to the target $Y$. In this way, the learned representations can be of good uniformity and sufficient predictive ability. Unlike IB that minimizes input-representation information, we handle the information in an opposite way to avoid the over-compression issue of latent representations.
@@ 2.1 Information Flow Maximization
As shown in Figure 1, regions being minimized by $I(X;Z)$ overlap with relevancy regions being maximized by $I(Y;Z)$. To preserve informative signals from the input $X$ relevant to the target $Y$ (i.e., region ①) as much as possible in the information flow of $X\rightarrow Z$, we use the idea of InfoMax Linsker (1988) to maximize $I(X;Z)$. Besides, we also maximize $I(Y;Z)$ to capture necessary information relevant to the target task $Y$. Combing the above two terms, the principle of information flow maximization (IFM) simultaneously maximizes both input-representation and representation-label mutual information. It can be formulated as the maximization of the following Lagrangian,
@@
In Equation (1), the second term that maximizing input-representation information $I(X;Z)$ ensures sufficiency of the representations for the input $X$ and preserves informative signals from $X$ relevant to the target $Y$. The first term that maximizing representation-label information $I(Y;Z)$ captures necessary information relevant to the target $Y$. In this way, the learned representation by IFM can be of good uniformity and sufficient predictive ability. Unlike the IB principle that minimizes input-representation information, we handle the information in an opposite way to avoid the over-compression issue of latent representations.
@@
The implementation of IFM contains two terms, i.e., maximizing $I(Y;Z)$ and $I(X;Z)$. First, we maximize the lower bound of $I(Y;Z)$ by estimating the conditional entropy of the target $Y$ given representations $Z$. Following Kolchinsky et al. (2019) and Hu et al. (2024), we use cross-entropy (CE) and mean squared error (MSE) as the estimators for classification and regression, respectively, i.e., $-\log\text{Softmax}(z_{i},y_{i})$ and $||z_{i}-y_{i}||^{2}$ for the $i$-th sample. Then, maximizing $I(X;Z)$ can be optimized by using the mutual information estimators, e.g., InfoNCE [^2] InfoNCE is a softmax-based version of noise-contrastive estimation Gutmann and Hyvärinen (2010). van den Oord et al. (2018) and MINE (mutual information neural estimator) Belghazi et al. (2018).
Here, we take InfoNCE as the default estimator for maximizing $I(X;Z)$. According to the information flow of $Z\leftarrow X\rightarrow Z^{\prime}$, the Markov chain rule states that $I(X;Z)\geq I(Z;Z^{\prime})$. Maximizing the lower bound of $I(X;Z)$, i.e., $I(Z;Z^{\prime})$, can be optimized by using InfoNCE estimator. Specifically, for a sample $x_{i}$ with the representation $z_{i}$, its positive key is the augmented sample obtained by dropout, and its negative keys are the other samples in the batch of $N$ examples. In this case, we have,
@@
where $z_{i}^{+}$ is the positive key and $z_{j}^{-}$ is the negative key. $K$ represents the number of negative keys, which is $N-1$ in the default mode. $s(\cdot)$ is a pairwise similarity function, i.e., cosine similarity, which can be seen as the dot product with $L_{2}$ normalization. $\tau>0$ is a scalar temperature parameter that controls the concentration or separation of probability distribution.
Alternatively, we also present how to utilize the MINE estimator to maximize $I(X;Z)$. Specifically, we use a neural network to approximate the lower bound of $I(X;Z)$, i.e.,
@@
where ${\mathbb{P}_{{X}{Z}}}$ is the joint distribution and ${\mathbb{P}_{X}\otimes\mathbb{P}_{Z}}$ is the product of the marginals. $T_{\omega}$ is a function parametrized by an MLP (i.e., a fully-connected neural network with two hidden layers) with the parameter $\omega\in\Omega$. The expectations of $I_{\Omega}(X;Z)$ are estimated by shuffling the samples from the joint distribution along the batch axis. Additionally, an exponential moving average is used to make an unbiased estimate.
@@ 2.2 Conditional Information Minimization
Inspired by Terzi et al. (2021), we use gradient-based adversarial training Goodfellow et al. (2015); Miyato et al. (2017) to approximately estimate the minimization of the Fisher information $F_{z|x,y}$ in the activations $Z_{\delta}$ given the input $X$ and target $Y$. Formally, denote $(x,y)$ as a mini-batch input sampled from distribution $D$ and $p(y|x;{\theta})$ as a model with the parameter $\theta$. The worst-case perturbation $\delta$ under the given $Y$ can be computed by the back-propagation in the network, i.e., $\max_{\|\delta\|_{q}\leq\epsilon}{{KL}(p(\cdot|x;{\theta_{\delta}});p(\cdot|x;{\theta}))}$ with a Lagrange constraint $\mathcal{L}_{IFM}(x,y;\theta_{\delta})$ and an $L_{q}$ norm constraint. At each step of training, we identify the adversarial perturbations $\delta$ against the current network with the parameter $\hat{\theta}$, and put them on the weights of first several hidden layers of the network. When using a Transformer-based network (e.g., RoBERTa) as the model backbone, we probabilistically put the perturbations on the weights of an embedding layer and the first hidden layer of the Transformer based on a predefined perturbation rate. With a linear approximation Goodfellow et al. (2015), an $L_{q}$ norm-ball constraint, a certain radius $\epsilon$ for $\delta$, and the original objective $\mathcal{L}_{IFM}$, the formulation of adversarial objective $\mathcal{L}_{CIM}$ is,
@@ 3.3 Ablation Study
Table 5: Transferability results (%) of GoEmotions $\rightarrow$ Others and SentiEval $\rightarrow$ Others with RoBERTa backbone. GoEmotions and SentiEval are the source datasets for pre-training. Pre-train Obj. indicates the pre-training objective. Cls. refers to the classifier when fine-tuning. Best results for linear (i.e., Linear) and nonlinear (i.e., CNN) classifiers are highlighted in bold. Detailed results of SentiEval $\rightarrow$ Others are listed in Appendix C.2.
@@ 3.4 Generalization Evaluation
We choose emotion-related benchmarks (e.g., EmotionEval, ISEAR, MELD, and GoEmotions), which are collected from different domains and aim to predict the label of different emotion taxonomies. To implement OOD scenarios, we train the model on the original training set from a source domain, select the best model based on the validation set of the source domain, and test on the test set of a target domain. To avoid the interference of label mapping bias between different taxonomies, each model is trained on the dataset with coarse-grained taxonomy to predict the label for another dataset with fine-grained taxonomy (i.e., GoEmotions). Table 4 shows the performance under OOD scenarios. CIFM obtains the best results on all OOD settings. Comparing to CE, CIFM achieves +7.9% improvements in terms of average scores. This results exhibit that CIFM can promote the learning of more sufficient representations for target task than others, which enhances the generalization in handling OOD scenarios across different domain shifts.
@@ 3.7 Representation Quality Evaluation
To evaluate the quality of the learned representations, we use two metrics to measure the sufficiency of both the input and target task. Following Wang and Isola (2020), we choose the uniformity metric to measure the preserved maximal information of hidden representations from the input. Following Hu et al. (2023), we apply the adjusted rand index (ARI) metric to assess the preserved maximal information of output representations for label structure. The lower the uniformity loss, the better sufficiency for the data. Conversely, a higher ARI score indicates a better sufficiency for the target.
@@ 4 Conclusion
We propose an information-theoretic CIFM framework to extract noise-invariant sufficient representations for the input data and target task. Firstly, an IFM principle is proposed to learn more sufficient representations for the input and target. Besides, a CIM principle is designed to eliminate negative redundant features while preserve noise-invariant features from the input. Experiments on 13 language understanding benchmarks demonstrate that CIFM effectively improves the performance of PLMs for classification and regression. Extensive experiments show that the learned representations are more sufficient, robust and transferable.
@@
This paper introduces a new information-theoretic framework to extract sufficient representations for the input data and target task, as well as eliminates negative redundant features from the input. It promotes learned representations with good uniformity and sufficient predictive ability, thereby enhancing the generalization of pre-trained language models for the target task. However, the technique is studied only for different classification and regression tasks; its effectiveness in generation tasks remains to be further validated. Additionally, the principles proposed in this paper require more in-depth theoretical investigations.
@@ Appendix A Related Work
Information-theoretic representation learning has become the workhorse of several impressive deep learning achievements, ranging from practical applications Pereyra et al. (2017); Alemi et al. (2017); Belghazi et al. (2018) to theoretical investigations Xu and Raginsky (2017); Steinke and Zakynthinou (2020); Kawaguchi et al. (2023). The aim is typically to discover useful and informative latent representations in a principled and systematic manner. In the context of the training principle, information-theoretic representation learning approaches can be broadly categorized into InfoMax-based and IB-based methods.
@@
In the field of machine learning, the information maximization (InfoMax) principle was first introduced by Linsker (1988). It suggests that a linear or nonlinear network can be viewed as an information channel and the goal is to transmit as much information as possible from the input data through the network. In recent years, InfoMax has extensive applications in the field of self-supervised representation learning Chen et al. (2016); van den Oord et al. (2018); Hjelm et al. (2019); Tschannen et al. (2020); Kong et al. (2020). Chen et al. (2016) extend generative adversarial networks to learn disentangled representations by similarly using InfoMax. Tschannen et al. (2020) study the usage of InfoMax in representation learning and shows the success depend on the inductive bias by the model architectures and mutual information (MI) estimators. Kong et al. (2020) explore several embedding models based on InfoMax, and introduce an objective that maximizes MI between the sentence representation and the n-grams. In supervised scenarios, minimizing the standard cross-entropy is actually equivalent to maximizing the mutual information between the representations and the target task Achille and Soatto (2018); Kolchinsky et al. (2019); Boudiaf et al. (2020). Moreover, Pereyra et al. (2017) add the negative entropy to the negative log-likelihood to penalize overconfident output distributions in a supervised manner.
@@
The information bottleneck (IB) principle Tishby et al. (1999); Tishby and Zaslavsky (2015) aims to balance the trade-off between the compression of representation and the power of predicting. Alemi et al. (2017) propose an efficient variational estimation method VIB based on the IB principle. Peng et al. (2019) use IB to constrain information flow in the discriminator for better adversarial learning. Then, Mahabadi et al. (2021) use VIB to suppress irrelevant features when fine-tuning on low-resource language understanding tasks. For tractable application of IB in a continuous setting, Belghazi et al. (2018) propose a mutual information neural estimation method to estimate MI and design an IB principle based on the estimator. Ragonesi et al. (2021) employ MINE-IB to learn unbiased representations. Fischer (2020) and Ramé and Cord (2021) introduce a conditional mutual information term to alleviate the over- and under-compression issue of IB. An et al. (2023) explicitly use the conditional entropy of the stochastic embedding as a confidence indicator and encourage the model to assign larger variance to more certain inputs. Recently, Hu et al. (2024) propose a probabilistic coding framework to mitigate the over-compression issue of IB by simultaneously performing compression and prediction using a shared learnable network. They also introduced a regularization term to promote class-level uniformity for better task prediction. Moreover, variational autoencoder (VAE) Kingma and Welling (2014) can be seen as a special case of an unsupervised VIB and have shown impressive results in self-supervised learning Sohl-Dickstein et al. (2015); Higgins et al. (2017); Hu et al. (2022).